Anna is now 3 years older than Sophie was a year ago. How old will Sophie be when Anna is 16?
Answer by stanbon(75887)

Anna is now 3 years older than Sophie was a year ago. How old will Sophie be when Anna is 16?
----
Equations:
a = (s-1)+3
---
Solve for "s":
s = a-2
---
s(16) = 16-2
---
s = 14 yrs (Sophie's age when anna is 16).
==================
Cheers,
Stan H.